Jane has a collection of 175 coins in her piggybank valued at $44.25 if she has 30 less dimes than Nicolas and 10 morehalf dollars than nickels , how many coins of each type ate in jane's piggybank? Answer by oberobic(2304) 
 With money problems you have to keep track of the number of coins and their values.
.
For example,
d = number of dimes
10d = value of the dimes in cents
n = number of nickels
5d = value of the nickels in cents
h = number of half dollars
50h = value of the half dollars in cents
.
n + d + h = 175 coins
5n + 10d + 50h = 4425 cents
.
d = n - 30
h = n + 10
.
substitute
.
n + n-30 + n+10 = 175
3n -20 = 175
3n = 195
n = 65
.
d = n-30
d = 65 -30
d = 35
.
h = n +10
h = 65 +10
h = 75
.
Always check your work.
In this case, what is the value of the coins?
5(65) = 325 cents
10(35) = 350 cents
50(75) = 3750 cents
325+350+3750= 4425 cents
Correct.
.
Answer: Jane has 65 nickels, 35 dimes, and 75 half dollars.
.
Done.